# Wet Hair: Global Symmetries in Entanglement Islands

### v1

@@ 3.2.3 Operators in the Island and Their Global Symmetry Charges
It is easy to check that the above operators Equ. (70) and Equ. (71) obey the constraint Equ. (69). [^6] We note that there is in fact a third way the dress operators to obey the constraint Equ. (69). This works only for special cases where we have oppositely charged operators. In this case, one can dress the oppositely charged operators using a Wilson line connecting them. We will not consider this possibility as the total global symmetry charge for the operators is zero. though the ways they obey the constraint are rather different. Equ. (70) obeys the constraint by dressing the original charged operator $\hat{\phi}(x,z)$ from $(x,z)$ to the boundary using a line operator and this dressing creates an operator from a charged operator situated at $(x,z)$ to a an operator with boundary $z=0$ charge. Hence the dressed operator Equ. (70) can be detected near the boundary by measuring its correlation functions with the near boundary gauge fields. This is the standard dressing in gauge theories.
@@
As was pointed out in Geng:2020qvw ; Geng:2021hlu ; Geng:2023zhq , a basic consistency condition for the holographic interpretation of entanglement islands is that the gravitational AdS operators inside the island have to commute with operators outside the island. This consistency condition enforces the second way of dressing for operators inside the island $\mathcal{I}$. Moreover, Geng:2025rov pointed out another basic consistency condition for the holographic interpretation of entanglement islands that operators inside the island have to be detectable by operators in the bath. For example, we can apply an operator in the island to create a local energetic excitation inside the island. Given the duality, the bath Hamiltonian should not commute with operators in the island.
@@ 3.3 Summary and Discussions
In summary, in this section we performed an in-depth study of symmetries in the island setup. We found that with leaky charged matter in the gravitational AdS part of the island setup, the gauge symmetry is spontaneously broken and is mixed with an exact global symmetry. The study of the gauge symmetry and its spontaneous breaking was performed in both the dual field description of the island setup and the island setup itself with various exactly matched results. This study extends the results and proved conjectures in various old and recent papers Porrati:2001gx ; Porrati:2002dt ; Porrati:2003sa ; Duff:2004wh ; Aharony:2006hz ; Rattazzi:2009ux ; Karch:2023wui with an explicit demonstration of the existence of a Goldstone boson, which is eaten by the gauge field in the Higgs mechanism, and the understanding of its field theory dual. The Goldstone boson is non-locally supported in the union of the AdS and the bath, though an observer inside the AdS will observe it as a local field. This fact enables us to identify operators inside the island that are consistent with the holographic interpretation of the island. Moreover, the Goldstone boson is non-linearly charged under various coupling-induced broken global symmetries including the one of the bath and is neutral under the exact global symmetry. This property ensures that the global symmetry charges of operators inside the island will not be changed due to the dressing and can be detected from the bath but are invisible to the boundary CFTd-1. This is consistent with the mechanism uncovered in Geng:2025rov for the information in the island being fully encoded in the bath. This mechanism is fully consistent with our analysis in Sec. 2 regarding the quantum mechanics of black holes and global symmetries. Moreover, our construction and analysis in this section can be easily generalized to non-Abelian gauge symmetries.
@@ 4 The Higher-Dimensional Holographic Dual– Karch-Randall Braneworld
In Sec. 3, we studied a model with an island in which the AdS and the bath are assumed to be weakly coupled with a double-trace deformation Equ. (5). We uncovered various important aspects of this model, including the existence of global symmetries and the mechanism for the global symmetry charges of operators inside the island to be fully encoded in the bath. We studied both the island theory directly and its field theory dual. Nevertheless, it was not very intuitive how the global symmetry charges inside the island are encoded in the bath. In this section, we will study the Karch-Randall braneworld Karch:2000ct ; Karch:2000gx ; Geng:2023qwm , which has a strongly coupled holographic dual island setup. We will see how the mechanism we uncovered in Sec. 3.1 is nicely geometrized to an intuitive picture in this framework. This picture will be helpful in understanding the resolution of the Harlow-Shagoulian paradox in Sec. 6.
@@ 4.1 The Karch-Randall Braneworld
One can also deduce the existence of the gravitational theory by applying the AdS/CFT correspondence to the bulk to dualize it to a gravitational theory on the brane coupled to a conformal field theory living on the leftover asymptotic boundary through a transparent boundary condition. This is consistent with the fact that the graviton on the brane has a mass, which can be seen by doing a KK analysis of the bulk graviton reduced from AdSd+1 to AdSd Karch:2000ct ; Geng:2025rov ; Geng:2025tba3 . We have already seen this result in the description of the theory with an island– the intermediate description– in which the AdSd graviton becomes massive due to bath coupling that induces spontaneous breaking of the diffeomorphism symmetry in the AdSd Porrati:2001gx ; Porrati:2002dt ; Porrati:2003sa ; Duff:2004wh ; Aharony:2006hz ; Geng:2023qwm ; Geng:2023ynk ; Geng:2025rov , analogously to the discussion of Sec. 3 for the gauge field. Thus the Karch-Randall braneworld is a holographic model for the island setup, where however the matter fields are strongly coupled as for any holographic theory. The dual island setup is often referred to as the intermediate description of the Karch-Randall braneworld Geng:2020qvw ; Geng:2020fxl ; Geng:2023qwm .
Consistent with the above interpretation, there are entanglement islands that can be found using holographic tools in this setup. The entanglement entropy of a bath subregion $R$ can be computed using the modified Ryu-Takayanagi (RT) formula Ryu:2006bv ; Ryu:2006ef ; Fujita:2011fp ; Takayanagi:2011zk ; Miao:2017gyt . As opposed to the RT formula in the standard AdS/CFT setup, the RT surface can now end on the Karch-Randall brane and one also has to minimize the surface area of the RT surface over its possible ending point on the brane Chen:2020uac ; Chen:2020hmv ; Geng:2020qvw ; Geng:2020fxl . This minimization over the brane ending point is actually the holographic dual of the minimization over possible island regions in the island formula Equ. (2). This minimization prescription is derived using replica wormholes Geng:2024xpj following the proof of the RT formula in the standard AdS/CFT correspondence Lewkowycz:2013nqa ; Dong:2013qoa ; Camps:2013zua ; Miao:2014nxa . Thus we have
@@ 4.2 Global and Gauge Symmetries in the Karch-Randall Braneworld
The holographic dual intermediate description can be decoded from the $(d+1)$-dimensional bulk by Kaluza-Klein (KK) decomposition. Thus, we will perform the KK decompositions for the bulk gauge field and the bulk gauge transform. The result is a tower of massive $d$-dimensional gauge fields and the corresponding gauge transforms. These gauge fields are the gauge fields in the holographic dual island setup. We will see that (as with the graviton Karch:2000ct ) there is no normalizable zero mode. [^9] In a future paper Geng:2026tba1 , we will argue that as with the graviton, there is an anomalously light mode that reduces to the zero mode in the zero curvature limit. Thus, all the $d$-dimensional gauge fields are massive. [^10] We note that this extra-dimensional Wilson line is distinct from the Wilson line along the brane directions as it looks like a local field from the perspective of the brane, rather than a nonlocal line operator.
Now we perform the explicit analysis. The upshot of these calculations will be that in the KR picture the dressing of operators via $\theta$ can be realized by a Wilson line traversing the bulk spacetime as depicted in Fig.11. Part of this system has been studied in Geng:2025rov , with which our discussion in the following overlaps. We will write the most general (d+1)-dimensional warped geometry in the bulk before we restrict to AdSd+1. We consider the bulk metric
@@
When the bulk is AdSd+1, the brane position $\rho_{B}$ is related to the brane tension $T$ as $T=(d-1)\tanh\rho_{B}$ Geng:2023qwm with the bulk $\rho$-coordinate $\rho\in[\rho_{B},\infty)$.
@@
In summary, in the dual island setup, we have a global symmetry that is supported on the bath and a tower of independent spontaneously broken U(1) gauge symmetries in the AdS. In the bulk, these symmetries assemble to a single U(1) gauge symmetry which is not spontaneously broken. Note that the matter fields in the dual island setup in the Karch-Randall setup are strongly coupled which is distinct from the simple toy model with only a single gauge boson we considered in Sec. 3.1. A scenario that is closer to the island setup we considered in Sec. 3 would have a gauge boson in the gravitational AdS part whose mass is entirely due to one-loop quantum corrections of the leaky matter fields. This corresponds more closely to the limit that the Karch-Randall brane is close to the asymptotic boundary, i.e. $\rho_{B}\rightarrow$-$\infty$. In this case, there is a distinct anomalously light KK mode, whose primary support is near the KR brane. This KK mode is the lightest normalizable mode $n=1$ of the bulk photon and its mass is scale separated from that of the other KK modes. We will further elaborate this point in a companion paper Geng:2026tba1 . To study the global symmetry in this model, we neglect the other KK modes so the truncated dual island setup is closer to the island setup we considered in Sec. 3.
@@ 4.4 Gravitating Bath
As it is discovered in Geng:2020fxl , in a deformed setup of the Karch-Randall braneworld, there is no entanglement island. In this setup, one has two Karch-Randall branes as shown in Fig. 12. In this case, the dual intermediate description is a modification of the island setup, where the bath is also gravitational. [^13] This is the key reason for the absence of entanglement islands Geng:2020fxl . Thus, in this case the tension between global symmetry and the Bekenstein-Hawking entropy bound persists and we don’t expect our construction of global symmetry charged operators inside the AdS would work in this case. To see that this is indeed the case, we notice that the zero mode of the KK wave equation Equ. (82) is now normalizable as the $\rho$ goes from the left to right brane in a finite range. Hence, in the intermediate description there is a massless gauge boson associated with this zero mode and we no longer have purely massive gauge bosons. In fact, in this case the intermediate theory is similar to the standard AdS/CFT scenario and the U(1) gauge symmetry associated with the massless gauge boson is dual to a global symmetry in a dual CFT. This dual CFT lives on the defect in Fig. 12. This dual CFT ensures the Bekenstein-Hawking entropy bound which forbids the existence of global symmetry in the intermediate description.
@@
Figure 9: A constant time slice of an AdSd+1 with a Karch-Randall brane. The green surface denotes the brane and it has AdSd geometry. All the dashed straight lines are in fact legitimate places for the brane to reside with the specific slice selected according to the brane tension Geng:2023qwm . The purple-shaded region behind the brane is cutoff. The red dot is the codimension one submanifold of the bulk asymptotic boundary (the thick black line) where it intersects the brane. This submanifold is also called the defect in the literature.
@@ 5 Explicit Examples from String Theory
The Karch-Randall setup can be engineered in string theory Karch:2000ct ; Karch:2001cw ; DHoker:2007hhe ; Aharony:2011yc ; Bachas:2018zmb ; Karch:2022rvr ; Huertas:2023syg . It consists of $D3$-branes in 10d Type IIB superstring theory ending on $D5$- and NS5-branes, preserving a quarter of the supersymmetries. [^14] That is we have eight supercharges preserved. The gravitational dual of this configuration is described by an asymptotic $AdS_{5}\times\mathbf{S}^{5}$ region (the bulk) connected to an $AdS_{4}\times\mathbf{M}^{6}$ geometry (which models the Karch-Randall brane) with $\mathbf{M}^{6}$ a manifold that smoothly caps off to the $\mathbf{S}^{5}$, see Fig.5.3. In the following subsections we discuss its gauge and gravitational nature.
@@ 5.1 The D3D3/$D5/NS5$ System in Type IIB String Theory
The setup that we are interested in is 4d $\mathcal{N}=4$ $SU(N)$ SYM in a semi-infinite spacetime with a 3d boundary defect, which hosts a 3d $\mathcal{N}=4$ SCFT. This boundary is considered in Gaiotto:2008ak , which can be described using the Hanany-Witten brane configurations Hanany:1996ie . The 3d system can be coupled to the 4d system in different ways with the 3d super(conformal)symmetries preserved, realizing different half-BPS boundary conditions of the 4d $\mathcal{N}=4$ SYM. Note that much of this section is review of the string construction. We talk about the relevant global symmetries in Sec. 5.3.
This field theory setup is realized in string theory as follows. We start with a stack of NN $D3$-branes extending along the directions 0123. The gauge theory living on this stack of $D3$-branes is the usual 4d ${\cal{N}}=4$ SYM $SU(N)$ theory. A boundary for this theory can be introduced by allowing the $D3$-branes to end on an appropriate configuration of NS5- and $D5$-branes, which are localized along the 3-direction and extend along 012 as well as additional transverse directions. To preserve the maximum supersymmetry, the NS5-branes are oriented along 012 456 and coincide with the $D3$-branes in the directions 789, while the $D5$-branes extend along 012 789 and share the same position as the $D3$-branes in the directions 456. All 5-branes are placed at the same location in the direction 3, thereby defining a boundary for the 4d $SU(N)$ theory. This boundary configuration is a Hanany-Witten Hanany:1996ie setup of $D3$-branes suspended in the 3-direction among NS5- and $D5$-branes, describing a 3d ${\cal{N}}=4$ CFT in the particular limit of all the 5-branes coinciding in the 3-direction.
As studied in Gaiotto:2008ak , how the $D3$-branes end on the NS5- and $D5$-branes is very constrained. The upshot is that the NS5-branes (also the $D5$s) get grouped into $n$ stacks, with each stack containing $n_{a}$ branes associated with an irreducible $SU(2)$ representation of dimension $K_{a}$ Gaiotto:2008ak . This means that each of the $n_{a}$ NS5-branes in the $a^{th}$ stack has $K_{a}$ $D3$-branes ending on it. The same applies to the $D5$-branes for which instead of $(n_{a},K_{a})$ we use $(m_{b},L_{b})$, leading to the constraint:

### v2

@@ 3.2.3 Operators in the Island and Their Global Symmetry Charges
It is easy to check that the above operators Equ. (70) and Equ. (71) obey the constraint Equ. (69). [^6] We note that there is in fact a third way the dress operators to obey the constraint Equ. (69). This works only for special cases where we have oppositely charged operators François and Ravera (2025). In this case, one can dress the oppositely charged operators using a Wilson line connecting them. We will not consider this possibility as the total global symmetry charge for the operators of this type is zero. though the ways they obey the constraint are rather different. Equ. (70) obeys the constraint by dressing the original charged operator $\hat{\phi}(x,z)$ from $(x,z)$ to the boundary using a line operator and this dressing creates an operator from a charged operator situated at $(x,z)$ to a an operator with boundary $z=0$ charge. Hence the dressed operator Equ. (70) can be detected near the boundary by measuring its correlation functions with the near boundary gauge fields. This is the standard dressing in gauge theories.
@@
As was pointed out in Geng and Karch (2020); Geng et al. (2021); Geng (2023a), a basic consistency condition for the holographic interpretation of entanglement islands is that the gravitational AdS operators inside the island have to commute with operators outside the island. This consistency condition enforces the second way of dressing for operators inside the island $\mathcal{I}$. Moreover, Geng (2025) pointed out another basic consistency condition for the holographic interpretation of entanglement islands that operators inside the island have to be detectable by operators in the bath. For example, we can apply an operator in the island to create a local energetic excitation inside the island. Given the duality, the bath Hamiltonian should not commute with operators in the island.
@@ 3.3 Summary and Discussions
In summary, in this section we performed an in-depth study of symmetries in the island setup. We found that with leaky charged matter in the gravitational AdS part of the island setup, the gauge symmetry is spontaneously broken and is mixed with an exact global symmetry. The study of the gauge symmetry and its spontaneous breaking was performed in both the dual field description of the island setup and the island setup itself with various exactly matched results. This study extends the results and proved conjectures in various old and recent papers Porrati (2002); Porrati and Starinets (2002); Porrati (2003); Duff et al. (2004); Aharony et al. (2006); Rattazzi and Redi (2009); Karch et al. (2024) with an explicit demonstration of the existence of a Goldstone boson, which is eaten by the gauge field in the Higgs mechanism, and the understanding of its field theory dual. The Goldstone boson is non-locally supported in the union of the AdS and the bath, though an observer inside the AdS will observe it as a local field. This fact enables us to identify operators inside the island that are consistent with the holographic interpretation of the island. Moreover, the Goldstone boson is non-linearly charged under various coupling-induced broken global symmetries including the one of the bath and is neutral under the exact global symmetry. This property ensures that the global symmetry charges of operators inside the island will not be changed due to the dressing and can be detected from the bath but are invisible to the boundary CFTd-1. This is consistent with the mechanism uncovered in Geng (2025) for the information in the island being fully encoded in the bath. This mechanism is fully consistent with our analysis in Sec. 2 regarding the quantum mechanics of black holes and global symmetries. Moreover, our construction and analysis in this section can be easily generalized to non-Abelian gauge symmetries.
@@ 4 The Higher-Dimensional Holographic Dual– Karch-Randall Braneworld
In Sec. 3, we studied a model with an island in which the AdS and the bath are assumed to be weakly coupled with a double-trace deformation Equ. (5). We uncovered various important aspects of this model, including the existence of global symmetries and the mechanism for the global symmetry charges of operators inside the island to be fully encoded in the bath. We studied both the island theory directly and its field theory dual. Nevertheless, it was not very intuitive how the global symmetry charges inside the island are encoded in the bath. In this section, we will study the Karch-Randall braneworld Karch and Randall (2001b, c); Geng (2023c), which has a strongly coupled holographic dual island setup. We will see how the mechanism we uncovered in Sec. 3.1 is nicely geometrized to an intuitive picture in this framework. This picture will be helpful in understanding the resolution of the Harlow-Shaghoulian paradox in Sec. 6.
@@ 4.1 The Karch-Randall Braneworld
One can also deduce the existence of the gravitational theory by applying the AdS/CFT correspondence to the bulk to dualize it to a gravitational theory on the brane coupled to a conformal field theory living on the leftover asymptotic boundary through a transparent boundary condition. This is consistent with the fact that the graviton on the brane has a mass, which can be seen by doing a KK analysis of the bulk graviton reduced from AdSd+1 to AdSd Karch and Randall (2001b); Geng (2025); Geng and Randall (2025). We have already seen this result in the description of the theory with an island– the intermediate description– in which the AdSd graviton becomes massive due to bath coupling that induces spontaneous breaking of the diffeomorphism symmetry in the AdSd Porrati (2002); Porrati and Starinets (2002); Porrati (2003); Duff et al. (2004); Aharony et al. (2006); Geng (2023c, b, 2025), analogously to the discussion of Sec. 3 for the gauge field. Thus the Karch-Randall braneworld is a holographic model for the island setup, where however the matter fields are strongly coupled as for any holographic theory. The dual island setup is often referred to as the intermediate description of the Karch-Randall braneworld Geng and Karch (2020); Geng et al. (2020); Geng (2023c).
Consistent with the above interpretation, there are entanglement islands that can be found using holographic tools in this setup. The entanglement entropy of a bath subregion $R$ can be computed using the modified Ryu-Takayanagi (RT) formula Ryu and Takayanagi (2006b, a); Fujita et al. (2011); Takayanagi (2011); Miao et al. (2017). As opposed to the RT formula in the standard AdS/CFT setup, the RT surface can now end on the Karch-Randall brane and one also has to minimize the surface area of the RT surface over its possible ending point on the brane Chen et al. (2020a, b); Geng and Karch (2020); Geng et al. (2020). This minimization over the brane ending point is actually the holographic dual of the minimization over possible island regions in the island formula Equ. (2). This minimization prescription is derived using replica wormholes Geng (2024) following the proof of the RT formula in the standard AdS/CFT correspondence Lewkowycz and Maldacena (2013); Dong (2014); Camps (2014); Miao and Guo (2015). Thus we have
@@ 4.2 Global and Gauge Symmetries in the Karch-Randall Braneworld
The holographic dual intermediate description can be decoded from the $(d+1)$-dimensional bulk by Kaluza-Klein (KK) decomposition. Thus, we will perform the KK decompositions for the bulk gauge field and the bulk gauge transform. The result is a tower of massive $d$-dimensional gauge fields and the corresponding gauge transforms. These gauge fields are the gauge fields in the holographic dual island setup. We will see that (as with the graviton Karch and Randall (2001b)) there is no normalizable zero mode. [^9] In a future paper Geng et al. (2026), we will argue that as with the graviton, there is an anomalously light mode that reduces to the zero mode in the zero curvature limit. Thus, all the $d$-dimensional gauge fields are massive. [^10] We note that this extra-dimensional Wilson line is distinct from the Wilson line along the brane directions as it looks like a local field from the perspective of the brane, rather than a nonlocal line operator.
Now we perform the explicit analysis. The upshot of these calculations will be that in the KR picture the dressing of operators via $\theta$ can be realized by a Wilson line traversing the bulk spacetime as depicted in Fig.11. Part of this system has been studied in Geng (2025), with which our discussion in the following overlaps. We will write the most general (d+1)-dimensional warped geometry in the bulk before we restrict to AdSd+1. We consider the bulk metric
@@
When the bulk is AdSd+1, the brane position $\rho_{B}$ is related to the brane tension $T$ as $T=(d-1)\tanh\rho_{B}$ Geng (2023c) with the bulk $\rho$-coordinate $\rho\in[\rho_{B},\infty)$.
@@
In summary, in the dual island setup, we have a global symmetry that is supported on the bath and a tower of independent spontaneously broken U(1) gauge symmetries in the AdS. In the bulk, these symmetries assemble to a single U(1) gauge symmetry which is not spontaneously broken. Note that the matter fields in the dual island setup in the Karch-Randall setup are strongly coupled which is distinct from the simple toy model with only a single gauge boson we considered in Sec. 3.1. A scenario that is closer to the island setup we considered in Sec. 3 would have a gauge boson in the gravitational AdS part whose mass is entirely due to one-loop quantum corrections of the leaky matter fields. This corresponds more closely to the limit that the Karch-Randall brane is close to the asymptotic boundary, i.e. $\rho_{B}\rightarrow$-$\infty$. In this case, there is a distinct anomalously light KK mode, whose primary support is near the KR brane. This KK mode is the lightest normalizable mode $n=1$ of the bulk photon and its mass is scale separated from that of the other KK modes. We will further elaborate this point in a companion paper Geng et al. (2026). To study the global symmetry in this model, we neglect the other KK modes so the truncated dual island setup is closer to the island setup we considered in Sec. 3.
@@ 4.4 Gravitating Bath
As it is discovered in Geng et al. (2020), in a deformed setup of the Karch-Randall braneworld, there is no entanglement island. In this setup, one has two Karch-Randall branes as shown in Fig. 12. In this case, the dual intermediate description is a modification of the island setup, where the bath is also gravitational. [^13] This is the key reason for the absence of entanglement islands Geng et al. (2020). Thus, in this case the tension between global symmetry and the Bekenstein-Hawking entropy bound persists and we don’t expect our construction of global symmetry charged operators inside the AdS would work in this case. To see that this is indeed the case, we notice that the zero mode of the KK wave equation Equ. (82) is now normalizable as the $\rho$ goes from the left to right brane in a finite range. Hence, in the intermediate description there is a massless gauge boson associated with this zero mode and we no longer have purely massive gauge bosons. In fact, in this case the intermediate theory is similar to the standard AdS/CFT scenario and the U(1) gauge symmetry associated with the massless gauge boson is dual to a global symmetry in a dual CFT. This dual CFT lives on the defect in Fig. 12. This dual CFT ensures the Bekenstein-Hawking entropy bound which forbids the existence of global symmetry in the intermediate description.
@@
Figure 9: A constant time slice of an AdSd+1 with a Karch-Randall brane. The green surface denotes the brane and it has AdSd geometry. All the dashed straight lines are in fact legitimate places for the brane to reside with the specific slice selected according to the brane tension Geng (2023c). The purple-shaded region behind the brane is cutoff. The red dot is the codimension one submanifold of the bulk asymptotic boundary (the thick black line) where it intersects the brane. This submanifold is also called the defect in the literature.
@@ 5 Explicit Examples from String Theory
The Karch-Randall setup can be engineered in string theory Karch and Randall (2001b, a); D’Hoker et al. (2007b); Aharony et al. (2011); Bachas and Lavdas (2018); Karch et al. (2022); Huertas and Uranga (2023). It consists of $D3$-branes in 10d Type IIB superstring theory ending on $D5$- and NS5-branes, preserving a quarter of the supersymmetries. [^14] That is we have eight supercharges preserved. The gravitational dual of this configuration is described by an asymptotic $AdS_{5}\times\mathbf{S}^{5}$ region (the bulk) connected to an $AdS_{4}\times\mathbf{M}^{6}$ geometry (which models the Karch-Randall brane) with $\mathbf{M}^{6}$ a manifold that smoothly caps off to the $\mathbf{S}^{5}$, see Fig.5.3. In the following subsections we discuss its gauge and gravitational nature.
@@ 5.1 The D3D3/$D5/NS5$ System in Type IIB String Theory
The setup that we are interested in is 4d $\mathcal{N}=4$ $SU(N)$ SYM in a semi-infinite spacetime with a 3d boundary defect, which hosts a 3d $\mathcal{N}=4$ SCFT. This boundary is considered in Gaiotto and Witten (2009), which can be described using the Hanany-Witten brane configurations Hanany and Witten (1997). The 3d system can be coupled to the 4d system in different ways with the 3d super(conformal)symmetries preserved, realizing different half-BPS boundary conditions of the 4d $\mathcal{N}=4$ SYM. Note that much of this section is review of the string construction. We talk about the relevant global symmetries in Sec. 5.3.
This field theory setup is realized in string theory as follows. We start with a stack of NN $D3$-branes extending along the directions 0123. The gauge theory living on this stack of $D3$-branes is the usual 4d ${\cal{N}}=4$ SYM $SU(N)$ theory. A boundary for this theory can be introduced by allowing the $D3$-branes to end on an appropriate configuration of NS5- and $D5$-branes, which are localized along the 3-direction and extend along 012 as well as additional transverse directions. To preserve the maximum supersymmetry, the NS5-branes are oriented along 012 456 and coincide with the $D3$-branes in the directions 789, while the $D5$-branes extend along 012 789 and share the same position as the $D3$-branes in the directions 456. All 5-branes are placed at the same location in the direction 3, thereby defining a boundary for the 4d $SU(N)$ theory. This boundary configuration is a Hanany-Witten Hanany and Witten (1997) setup of $D3$-branes suspended in the 3-direction among NS5- and $D5$-branes, describing a 3d ${\cal{N}}=4$ CFT in the particular limit of all the 5-branes coinciding in the 3-direction.
As studied in Gaiotto and Witten (2009), how the $D3$-branes end on the NS5- and $D5$-branes is very constrained. The upshot is that the NS5-branes (also the $D5$s) get grouped into $n$ stacks, with each stack containing $n_{a}$ branes associated with an irreducible $SU(2)$ representation of dimension $K_{a}$ Gaiotto and Witten (2009). This means that each of the $n_{a}$ NS5-branes in the $a^{th}$ stack has $K_{a}$ $D3$-branes ending on it. The same applies to the $D5$-branes for which instead of $(n_{a},K_{a})$ we use $(m_{b},L_{b})$, leading to the constraint: